Question 16, chap 126, sect 1. part I of 1 10 points A capacitor that is connected to a 45 V
source contains 88 uC of charge. What is its capacitance? 1. 1.83673 μ 、 2. 1.89583 μ!.. 3. 1.95556 μF 4.2 02011 μF 5. 2.08333 p
Answer #1

given

voltage v = 45 V

charge q = 88*10^-6 C

from the relation

charge q = cv

capacitance c = q/v

c = 88*10^-6/45 = 1.95556*10^-6 F

option 3